OpenStudy (anonymous):

(-3 - 2)^2 - 30 divided by 5 * 2 + 4^2

OpenStudy (anonymous):

25 -30 = -5 / 10+8 = -5/18

OpenStudy (anonymous):

Oh no it was just like 30 divided by 5 not whole equation before the 30 I just didn't have a divide sign